# Gauri Pradhan

### v1

@@ 8.1 Importance of Choosing Fitting Hyperparameters for LiRA:
Using good hyperparameters to train the shadow models is critical to the performance of LiRA. As observed in Section 5, ACC-LiRA performs considerably worse than LiRA and KL-LiRA, especially at low fpr. It can be seen from Figure A.20 that the variance of the IN and OUT distributions for data samples is much larger for ACC-LiRA as compared to LiRA or KL-LiRA. ACC-LiRA differs from the other two attacks in its approach to choose optimal hyperparameters to train the shadow models. We suspect that the large variance of the distributions is due to each shadow model being trained with a different set of hyperparameters with the aim to maximize their utility on the corresponding shadow data sets.
Additionally, Figure A.21 shows that the hyperparameters (learning rate and batch size) found using ACC-LiRA are not close to the target hyperparameters. Whereas for KL-LiRA, the learning rate used for training the shadow models is similar to the target learning rate. This makes it difficult for ACC-LiRA to distinguish between the 2 distributions thus inhibiting its ability to label the data samples as IN or OUT with confidence.
@@ 9 Conclusion
In this paper, we demonstrated that score-based MIAs that use shadow models (e.g. LiRA) do not require prior knowledge of target model’s hyperparameters as a necessary pre-condition for success, provided that an alternate set of hyperparameters is chosen to train the shadow models to optimize the efficiency of the attack. To facilitate this, we introduced KL-LiRA which selects hyperparameters for shadow model training based on similarity between the output distributions of the target model and shadow models.
In our experiments, KL-LiRA’s performance was very similar to an attack using target model’s hyperparameters against non-DP models, thereby showing that keeping the hyperparameters secret provides no additional privacy protection for a non-DP model. For models trained with DP, KL-LiRA performed slightly worse than an attack using target model’s hyperparameters.
We also studied whether using training data for HPO under DP leads to privacy leakage that is significant enough to be detected by an empirical privacy attack such as MIA. We found no evidence that suggests that using training data for HPO would lead to significant additional privacy leakage in the transfer learning setting.
@@ A.1 Hyperparameter Optimization (HPO)
Fig. A.12: MIA efficiency (tpr) against WR-50 + Head target models when compared to MIA efficiency against R-50 + Head target models trained on CIFAR-10 ($S=50$). WR-50 is trained on ImageNet-1K making it a weaker pretrained model than R-50 trained on ImageNet-21K. Performance of all versions of LiRA improve with a weaker pretrained backbone implying this to be an easy setting for attacks. For each of these attacks, we use $M=128$ shadow models per target model. Each target model is trained on CIFAR10 ($S=50$) data set.
@@ A.2 Attacking Models Using Weaker Pretrained Backbone
To evaluate the effect of using a weaker pretrained backbone for the target model, we fine-tune Wide-ResNet-50 (WR-50) with $68.9$M parameters [56] pretrained on ImageNet-1K [39] using Head as parameterization as our target models for 120 training epochs. In our experiment, we consider the $\texttt{KNOWN}(\mathcal{A})$ setting. Therefore, the shadow models have the same architecture as the target models. The results are shown in Figure A.12. Using a pretrained model trained on a subset of ImageNet-21K tends to enhance the attack efficiency of all 3 variants of LiRA indicating that this is an easy setting for attacks. Evidently, the performance of KL-LiRA relative to LiRA suffers when a weaker pretrained model is used for transfer learning but not significantly. tpr for KL-LiRA is 0.85×0.85$\times$0.85 × of the tpr for LiRA at $0.1\%$ fpr with WR-50 while it is 0.88×0.$8$ of the tpr for LiRA at $0.1\%$ fpr with R-50 [21].
Fig. A.13: Evaluating attack performance against models trained from scratch using MNIST data set calculated over $100$ target models. Each target model was attacked using $99$ shadow models. Similar to deep transfer learning setting, we find that KL-LiRA is the optimal choice for attack when the attacker has no knowledge of the target model’s hyperparameters.
@@ A.3 Attacking Models Trained From Scratch
To evaluate the efficiency of KL-LiRA beyond the deep transfer learning setting, we train a CNN with a convolutional layer (with 16 filters) followed by a fully connected layer (with 10 units) from scratch using MNIST data set [25]. We train 100 such models as the target models for the attack. For training from scratch we run 100 HPO trials using Optuna for each target model. Our preliminary analysis as shown in Figure A.13 demonstrates that KL-LiRA manages to match the attack success rate (tpr) of LiRA at low fprs and is a superior choice of attack to ACC-LiRA in this setting.
@@ A.4 Varying the number of unknown hyperparameters in HPO
In this section, we briefly explores how KL-LiRA fairs when we expand the set of hyperparameters unknown to the attacker to include the exact number of epochs used for training the target model. During HPO, we maintain a range of $\{1,200\}$ for the number of training epochs. From Figure A.14, it can be seen that KL-LiRA continues to be the better approach to find optimal hyperparameters for training the shadow models as the pool of hyperparameters not known to the attack expands to include the number of epochs used to train the target model. The power of KL-LiRA suffers with a larger set of unknown hyperparameters which requires further investigation in the future. tpr for KL-LiRA (when tuning for training epochs) is 0.62×0.62$\times$0.62 × lower than tpr for KL-LiRA (without tuning for training epochs) at $0.1\%$ fpr.
Fig. A.14: Impact of adding epochs to the set of hyperparameters unknown to the attack in $\texttt{KNOWN}(\mathcal{A})$ threat model on the MIA efficiency. Increasing the set of unknown hyperparameters decreases the efficiency of KL-LiRA with respect to LiRA. For comparison, we also depict the performance of different attacks with fixed epochs ($=40$) as baseline. In the latter setting, the optimizer hyperparameters unknown to the attacker include learning rate and batch size. For each of these attacks, we use $M=128$ shadow models per target model. Each target model is trained on CIFAR10 ($S=50$) with ViT-B + Head architecture.
@@ A.5 Additional Figures and Tables
Fig. A.15: Attack performances against models protected by DP ($\varepsilon=8,\delta=10^{-5}$). ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. Using DP has adverse effect on the attacks’ success rate. Under DP, WB-LiRA is the most powerful attack followed by KL-LiRA and ACC-LiRA.
@@
Fig. A.16: Attack performances against models protected by DP ($\varepsilon=1,\delta=10^{-5}$). ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. A moderate privacy level of $\varepsilon=1$ reduces the power of the attacks close to random guessing.
@@
Fig. A.17: The plot compares the average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments grouped by privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
@@
Fig. A.18: The plot compares the average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments and privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr on a logarithmic scale. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
@@
Fig. A.19: The plot above shows $p$-values (unadjusted and adjusted by Benjamini-Yekutieli (BY) procedure) on a negative logarithmic scale obtained for the permutation tests comparing the distribution of tpr at fixed $\textsc{fpr}(0.001,0.01,0.1)$ for TD-HPO (where HPO is done using the training data), $\mathbb{P}_{\text{TD}}(\textsc{tpr}|\textsc{fpr})$, and ED-HPO (where HPO is done using external data), $\mathbb{P}_{\text{ED}}(\textsc{tpr}|\textsc{fpr})$, for all the experiments sorted in increasing order. The negative-logarithmic scale highlights the $p$-values that are significant. Additionally, we provide the details of the experiments for which the $p$-value is above the significance level or stands out among the adjusted $p$-values.
Fig. A.20: Samplewise IN and OUT score distributions for different attacks for samples drawn from CIFAR-100. The IN and OUT distributions for ACC-LiRA happen to show considerable overlapping compared to the corresponding score distributions for WB-LiRA or KL-LiRA.
@@
Fig. A.21: Scatter plots depicted relation between the hyperparameters used for LiRA (target model’s hyperparameters) and the hyperparameters found using ACC-LiRA and KL-LiRA to train the shadow models. Learning rate found using KL-LiRA is close to the learning rate of the target model. For ACC-LiRA, neither the learning rate nor the batch size seems to have any meaningful relationship with the target model’s hyperparameters as observed by the lack of a definite pattern in the corresponding scatter plots. Same is true for the relationship between batch size of the shadow models and the target models in KL-LiRA setting.

### v2

@@ 8.1 Importance of Choosing Fitting Hyperparameters for LiRA:
Using good hyperparameters to train the shadow models is critical to the performance of LiRA. As observed in Section 5, ACC-LiRA performs considerably worse than LiRA and KL-LiRA, especially at low fpr. It can be seen from Figure A9 that the variance of the IN and OUT distributions for data samples is much larger for ACC-LiRA as compared to LiRA or KL-LiRA. ACC-LiRA differs from the other two attacks in its approach to choose optimal hyperparameters to train the shadow models. We suspect that the large variance of the distributions is due to each shadow model being trained with a different set of hyperparameters with the aim to maximize their utility on the corresponding shadow data sets.
Additionally, Figure A10 shows that the hyperparameters (learning rate and batch size) found using ACC-LiRA are not close to the target hyperparameters. Whereas for KL-LiRA, the learning rate used for training the shadow models is similar to the target learning rate. This makes it difficult for ACC-LiRA to distinguish between the 2 distributions thus inhibiting its ability to label the data samples as IN or OUT with confidence.
@@ 9 Conclusion
In this paper, we demonstrated that score-based MIAs that use shadow models (e.g. LiRA) do not require prior knowledge of target model’s hyperparameters as a necessary pre-condition for success, provided that an alternate set of hyperparameters is chosen to train the shadow models to optimize the efficiency of the attack. To facilitate this, we introduced KL-LiRA which selects hyperparameters for shadow model training based on similarity between the output distributions of the target model and shadow models. In our experiments, KL-LiRA’s performance was very similar to an attack using target model’s hyperparameters against non-DP models, thereby showing that keeping the hyperparameters secret provides no additional privacy protection for a non-DP model. For models trained with DP, KL-LiRA performed slightly worse than an attack using target model’s hyperparameters. We also studied whether using training data for HPO under DP leads to privacy leakage that is significant enough to be detected by an empirical privacy attack such as MIA. We found no evidence that suggests that using training data for HPO would lead to significant additional privacy leakage in the transfer learning setting.
@@ A.1 Hyperparameter Optimization
Fig. A1: MIA efficiency (tpr) against WR-50 + Head target models when compared to MIA efficiency against R-50 + Head target models trained on CIFAR-10 ($S=50$). WR-50 is trained on ImageNet-1K making it a weaker pretrained model than R-50 trained on ImageNet-21K. Performance of all versions of LiRA improve with a weaker pretrained backbone implying this to be an easy setting for attacks. For each of these attacks, we use $M=128$ shadow models per target model. Each target model is trained on CIFAR10 ($S=50$) data set.
Fig. A2: Evaluating attack performance against models trained from scratch using MNIST data set calculated over $100$ target models. Each target model was attacked using $99$ shadow models. Similar to deep transfer learning setting, we find that KL-LiRA is the optimal choice for attack when the attacker has no knowledge of the target model’s hyperparameters.
@@ A.2 Attacking Models Using Weaker Pretrained Backbone
To evaluate the effect of using a weaker pretrained backbone for the target model, we fine-tune Wide-ResNet-50 (WR-50) with $68.9$M parameters [56] pretrained on ImageNet-1K [38] using Head as parameterization as our target models for 120 training epochs. In our experiment, we consider the $\texttt{KNOWN}(\mathcal{A})$ setting. Therefore, the shadow models have the same architecture as the target models. The results are shown in Figure A1. Using a pretrained model trained on a subset of ImageNet-21K tends to enhance the attack efficiency of all 3 variants of LiRA indicating that this is an easy setting for attacks. Evidently, the performance of KL-LiRA relative to LiRA suffers when a weaker pretrained model is used for transfer learning but not significantly. tpr for KL-LiRA is 0.85×0.85$\times$0.85 × of the tpr for LiRA at $0.1\%$ fpr with WR-50 while it is 0.88×0.$8$ of the tpr for LiRA at $0.1\%$ fpr with R-50 [39].
@@ A.3 Attacking Models Trained From Scratch
To evaluate the efficiency of KL-LiRA beyond the deep transfer learning setting, we train a CNN with a convolutional layer (with 16 filters) followed by a fully connected layer (with 10 units) from scratch using MNIST data set [57]. We train 100 such models as the target models for the attack. For training from scratch we run 100 HPO trials using Optuna for each target model. Our preliminary analysis as shown in Figure A2 demonstrates that KL-LiRA manages to match the attack success rate (tpr) of LiRA at low fprs and is a superior choice of attack to ACC-LiRA in this setting.
Fig. A3: Impact of adding epochs to the set of hyperparameters unknown to the attack in $\texttt{KNOWN}(\mathcal{A})$ threat model on the MIA efficiency. Increasing the set of unknown hyperparameters decreases the efficiency of KL-LiRA with respect to LiRA. For comparison, we also depict the performance of different attacks with fixed epochs ($=40$) as baseline. In the latter setting, the optimizer hyperparameters unknown to the attacker include learning rate and batch size. For each of these attacks, we use $M=128$ shadow models per target model. Each target model is trained on CIFAR10 ($S=50$) with ViT-B + Head architecture.
@@ A.4 Varying the Number of Unknown Hyperparameters in HPO
In this section, we briefly explores how KL-LiRA fairs when we expand the set of hyperparameters unknown to the attacker to include the exact number of epochs used for training the target model. During HPO, we maintain a range of $\{1,200\}$ for the number of training epochs. From Figure A3, it can be seen that KL-LiRA continues to be the better approach to find optimal hyperparameters for training the shadow models as the pool of hyperparameters not known to the attack expands to include the number of epochs used to train the target model. The power of KL-LiRA suffers with a larger set of unknown hyperparameters which requires further investigation in the future. tpr for KL-LiRA (when tuning for training epochs) is 0.62×0.62$\times$0.62 × lower than tpr for KL-LiRA (without tuning for training epochs) at $0.1\%$ fpr.
@@ A.5 Additional Figures and Tables
Fig. A4: Attack performances against models protected by DP ($\varepsilon=8,\delta=10^{-5}$). ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. Using DP has adverse effect on the attacks’ success rate. Under DP, LiRA is the most powerful attack followed by KL-LiRA and ACC-LiRA.
@@
Fig. A5: Attack performances against models protected by DP ($\varepsilon=1,\delta=10^{-5}$). ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. A moderate privacy level of $\varepsilon=1$ reduces the power of the attacks close to random guessing.
@@
Fig. A6: The plot compares average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments grouped by privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
@@
Fig. A7: The plot compares average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments and privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr on a logarithmic scale. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
@@
Fig. A8: The plot above shows $p$-values (unadjusted and adjusted by Benjamini-Yekutieli (BY) procedure) on a negative logarithmic scale obtained for the permutation tests comparing the distribution of tpr at fixed $\textsc{fpr}(0.001,0.01,0.1)$ for TD-HPO (where HPO is done using the training data), $\mathbb{P}_{\text{TD}}(\textsc{tpr}|\textsc{fpr})$, and ED-HPO (where HPO is done using external data), $\mathbb{P}_{\text{ED}}(\textsc{tpr}|\textsc{fpr})$, for all the experiments sorted in increasing order. The negative-logarithmic scale highlights the $p$-values that are significant. Additionally, we provide the details of the experiments for which the $p$-value is above the significance level or stands out among the adjusted $p$-values.
Fig. A9: Samplewise IN and OUT score distributions for different attacks for samples drawn from CIFAR-100. The IN and OUT distributions for ACC-LiRA happen to show considerable overlapping compared to the corresponding score distributions for LiRA or KL-LiRA.
@@
Fig. A10: Scatter plots depicting relation between the hyperparameters used for LiRA (target model’s hyperparameters) and the hyperparameters found using ACC-LiRA and KL-LiRA to train the shadow models. Learning rate found using KL-LiRA is close to the learning rate of the target model. For ACC-LiRA, neither the learning rate nor the batch size seems to have any meaningful relationship with the target model’s hyperparameters as observed by the lack of a definite pattern in the corresponding scatter plots. Same is true for the relationship between batch size of the shadow models and the target models in KL-LiRA setting.